OpenStudy (anonymous):

find the solution to the system of equations \[3b-a=-7\] \[5a+6b=14\]

hero (hero):

Multiply both sides of the first equation by 5 15b - 5a = -35 5a + 6b = 14 Then add both equations together: (5a - 5a) + (15b + 6b) = (14 - 35) Then finish solving for b. Afterwards, solve for a

OpenStudy (anonymous):

got it! thank you!! :) i couldn't remember the rule :)